# Gaussian Head Avatar: Ultra High-fidelity Head Avatar via Dynamic Gaussians

### v2

@@ Abstract
Creating high-fidelity 3D head avatars has always been a research hotspot, but there remains a great challenge under lightweight sparse view setups. In this paper, we propose Gaussian Head Avatar represented by controllable 3D Gaussians for high-fidelity head avatar modeling. We optimize the neutral 3D Gaussians and a fully learned MLP-based deformation field to capture complex expressions. The two parts benefit each other, thereby our method can model fine-grained dynamic details while ensuring expression accuracy. Furthermore, we devise a well-designed geometry-guided initialization strategy based on implicit SDF and Deep Marching Tetrahedra for the stability and convergence of the training procedure. Experiments show our approach outperforms other state-of-the-art sparse-view methods, achieving ultra high-fidelity rendering quality at 2K resolution even under exaggerated expressions. Project page: https://yuelangx.github.io/gaussianheadavatar.
Figure 1: Gaussian head avatar achieves ultra high-fidelity image synthesis with controllable expressions at 2K resolution. The above shows different views of the synthesized avatar, and the bottom shows different identities animated by the same expression. 16 views are used during the training.
@@ 1 Introduction
High-fidelity 3D human head avatar modeling is of great significance in many fields, such as VR/AR, telepresence, digital human and film production. Automatically creating high-fidelity avatars has been a research hotspot in computer vision for decades. Although some traditional head avatars [37, 39, 41, 54] can realize high-fidelity animation, they typically require accurate geometries reconstructed and tracked from dense multi-view videos, thus limiting their applications in lightweight settings.
@@
Thanks to the Neural Radiance Fields (NeRF) [43] which show great capability of novel view synthesis in the absence of accurate geometry, recent methods [38, 57] skip the geometry reconstruction and tracking steps but directly learn high-quality NeRF-based head avatars. Other works [71, 48, 42] have verified that NeRF can be applied to either dense or sparse views, which greatly lowers the threshold for head avatar reconstruction. However, it still remains challenging for these NeRF-based approaches to synthesize high-fidelity images at 2K resolutions with pixel-level details, including wrinkles and eyes.
To overcome this bottleneck and further improve the avatar quality, we introduce 3D Gaussian splatting [26] for 3D head avatar modeling. This is an explicit discrete representation that can be well adapted to the rasterization-based rendering pipeline. It has been verified that the 3D Gaussian representation is capable of rendering complex scenes with low computational consumption. Compared to NeRF, the reconstruction quality of static and dynamic scenes [59, 40, 67] is much better while rendering time cost has been significantly reduced. Motivated by this progress, we propose Gaussian Head Avatar, a novel representation that utilizes 3D Gaussian splatting for ultra high-fidelity head avatar modeling. Although recent 4D Gaussian works [59, 67, 40, 66] have been proposed to reconstruct dynamic scenes, all of them cannot be animated. For modeling the animatable head avatar, it is crucial but still unexplored how to effectively control the deformation of 3D Gaussians and model the dynamic appearances through expression coefficients.
@@
Previous explicit [74] and implicit [2, 73, 76] head avatars usually formulate the facial deformation via linear blend skinning (LBS) using the skinning weights and blendshapes like the FLAME model [32]. However, such an LBS-based formulation fails to represent exaggerated and fine-grained expressions by simple linear operations, limiting the representation ability of the head avatars. Inspired by NeRSemble [28], we propose a fully learnable expression-conditioned deformation field for the 3D head Gaussians, avoiding the limited capability of the LBS-based formulation. Specifically, we input the positions of the 3D Gaussians with expression coefficients into an MLP to directly predict the displacements from the neutral expression to the target one. Similarly, we control the motion of non-face areas, such as the neck, using the head pose as the condition. 3D Gaussian-based representation has the powerful ability to reconstruct high-frequency details, enabling our method to learn accurate deformation fields. In turn, the learned accurate deformation field facilitates the dynamic Gaussian head model to fit more dynamic details. As a result, our method is able to reconstruct finer-grained dynamic details of expressive human heads.
@@
Unfortunately, as a discretized representation, the gradients back-propagated to the 3D Gaussians cannot spread through the whole space. Thus the convergence of training heavily relies on a plausible initialization for both the geometry and the deformation field. However, simply initializing the 3D head Gaussians with a morphable template like FLAME [32] fails to model the long hairstyle and the shoulders. Hence, we further propose an efficient and well-designed geometry-guided initialization strategy. Specifically, instead of starting from stochastic Gaussians or a FLAME model, we initially optimize an implicit signed distance function (SDF) field along with a color field and a deformation MLP for modeling the basic geometry, color, and the expression-conditioned deformations of the head avatar respectively. The SDF field is converted to a mesh through Deep Marching Tetrahedra (DMTet) [52], with the color and deformation of the vertices predicted by the MLPs. Then we render the mesh and optimize them jointly under the supervision of multi-view RGB images. Finally, we use the mesh with per-vertex features from the SDF field to initialize the 3D Gaussians to lie on the basic head surface while the color and deformation MLPs are carried over to the next stage, ensuring stable training for convergence. The entire initialization process takes only around 10 minutes.
The contributions of our method can be summarized as:
  • We propose Gaussian Head Avatar, a new head avatar representation that employs controllable dynamic 3D Gaussians to model expressive human head avatars, producing ultra high-fidelity synthesized images at 2K resolutions.
@@
  • For modeling high-frequency dynamic details, we employ a fully learned deformation field upon the 3D head Gaussians, which accurately model extremely complex and exaggerated facial expressions.
@@
  • We carefully design an efficient initialization strategy that leverages implicit representations to initialize the geometry and deformation, leading to efficient and robust convergence when training the Gaussian Head Avatar.
@@
Benefiting from these contributions, our method surpasses recent state-of-the-art methods under lightweight sparse-view setups on the avatar quality by a large margin.
2 Related Works
@@
3D Head Avatar Reconstruction. Due to the wide application value in the film and digital human industry, 3D head avatar reconstruction from multi-view images has always been a research hotspot. Traditional works [31, 4, 19, 8] reconstruct the scan geometry through multi-view stereo and then register a face mesh template to it. However, such methods usually require heavy computation. With the utilization of deep neural networks, current methods [33, 7, 61, 65] achieve very fast reconstruction, producing even more accurate geometry. Lombardi et al. [37], Bi et al. [5] and Ma et al. [41] represent the full head mesh through a deep neural network and train it with multi-view videos as supervision. However, due to the errors in geometric estimation, mesh-based head avatars typically suffer from texture blur. Therefore, some recent methods [57, 38] utilize NeRF representation [43] to synthesize novel view images without geometry reconstruction, or build NeRF on the head mesh template [39]. Furthermore, the NeRF-based methods are extended to sparse view reconstruction tasks [71, 42, 48, 28] and achieve impressive performance.
@@
Point-based Rendering. Point elements as a discrete and unstructured representation can fit geometry with arbitrary topology [69] efficiently. Recent methods [58, 30, 29] open up a differentiable rasterization pipeline, such that the point-based representation is widely used in multi-view reconstruction tasks. Aliev et al. [1] and Ruckert et al. [49] propose to first render the feature map, which is transferred to the images through a convolutional renderer. Xu et al.[62] use neural point cloud associated with neural features to model a NeRF. Recently, 3D Gaussian splatting [26] shows its superior performance, beating NeRF in both novel view synthesis quality and rendering speed. Some approaches [59, 67, 40, 66, 34, 72, 72, 50] extend Gaussian representation to dynamic scene reconstruction. However, these methods can not be migrated to the head avatar reconstruction tasks.
3 Overview
Figure 2: The overview of the Gaussian Head Avatar rendering and reconstruction. We first optimize the guidance model including a neutral mesh, a deformation MLP and a color MLP in the Initialization stage. Then we use them to initialize the neutral Gaussians and the dynamic generator. Finally, 2K RGB images are synthesized through differentiable rendering and the super-resolution network. The Gaussian Head Avatar are trained under the supervision of multi-view RGB videos.
The pipeline of the reconstruction of Gaussian Head Avatar is illustrated in Fig. 2, including the initialization stage and the training stage of Gaussian Head Avatar. Before the beginning of the pipeline, we remove the background [35] of each image and jointly estimate the 3DMM model [18], 3D facial landmarks and the expression coefficients for each frame. In the initialization stage (Sec. 4.3), we reconstruct an SDF-based neutral geometry, and optimize a deformation MLP and a color MLP from the training data as the guidance model. Next, we extract the neutral mesh through DMTet to initialize the neutral Gaussians while the deformation and color MLPs are also inherited from the initialization stage. In the training stage of Gaussian Head Avatar (Sec. 4.2), we deform the neutral Gaussians to the target expression through the dynamic generator given the driving expression coefficients as the condition. Finally, given a camera view, the expressive Gaussians are rendered to a feature map, which is then fed into the convolutional super resolution network to generate high-resolution avatar images. The whole model is optimized under the supervision of multi-view RGB videos.
@@ 4.1 Avatar Representation
Our task is to reconstruct a dynamic head avatar controlled by expression coefficients. Therefore, we formulate the head avatar as dynamic 3D Gaussians conditioned on expressions. To handle the dynamic changes, we input the expression coefficients with head pose to the head avatar model and output the position and other attributes of the Gaussians as above.
@@
Specifically, we first construct a canonical neutral Gaussian model with expression-independent attributes: $\{\boldsymbol{X}_{0},\boldsymbol{F}_{0},\boldsymbol{Q}_{0},\boldsymbol{S}_{0},\boldsymbol{A}_{0}\}$, which are fully optimizable. $\boldsymbol{X}_{0}\in\mathbb{R}^{N\times 3}$ denotes the positions of the Gaussians with a neutral expression in the canonical space. $\boldsymbol{F}_{0}\in\mathbb{R}^{N\times 128}$ denotes the point-wise feature vectors as their intrinsic properties. $\boldsymbol{Q}_{0}\in\mathbb{R}^{N\times 4}$, $\boldsymbol{S}_{0}\in\mathbb{R}^{N\times 3}$ and $\boldsymbol{A}_{0}\in\mathbb{R}^{N\times 1}$ denotes the neutral rotation, scale and opacity respectively. Note that we do not define the neutral color, but directly predict expression-dependent dynamic color from the point-wise feature vectors $\boldsymbol{F}_{0}$. Then, we construct an MLP-based expression conditioned dynamic generator $\boldsymbol{\Phi}$ to generate all the extra dynamic changes to the neutral model. Overall, the whole Gaussian head avatar can be formulated as:
$\{X,C,Q,S,A\}=\boldsymbol{\Phi}(\boldsymbol{X}_{0},\boldsymbol{F}_{0},\boldsymbol{Q}_{0},\boldsymbol{S}_{0},\boldsymbol{A}_{0};\theta,\beta),$ (2)
@@
with $\theta$ denoting expression coefficients and $\beta$ denoting the head pose. During the training, we optimize all the parameters of the dynamic generator $\boldsymbol{\Phi}$ and the neutral Gaussian model $\{\boldsymbol{X}_{0},\boldsymbol{F}_{0},\boldsymbol{Q}_{0},\boldsymbol{S}_{0},\boldsymbol{A}_{0}\}$, which are highlighted in bold in the following.

### v3

# HHAvatar: Gaussian Head Avatar with Dynamic Hairs

@@ Abstract
Creating high-fidelity 3D head avatars has always been a research hotspot, but it remains a great challenge under lightweight sparse view setups. In this paper, we propose HHAvatar represented by controllable 3D Gaussians for high-fidelity head avatar with dynamic hair modeling. We first use 3D Gaussians to represent the appearance of the head, and then jointly optimize neutral 3D Gaussians and a fully learned MLP-based deformation field to capture complex expressions. The two parts benefit each other, thereby our method can model fine-grained dynamic details while ensuring expression accuracy. Furthermore, we devise a well-designed geometry-guided initialization strategy based on implicit SDF and Deep Marching Tetrahedra for the stability and convergence of the training procedure. To address the problem of dynamic hair modeling, we introduce a hybrid head model into our avatar representation based Gaussian Head Avatar and a training method that considers timing information and an occlusion perception module to model the non-rigid motion of hair. Experiments show that our approach outperforms other state-of-the-art sparse-view methods, achieving ultra high-fidelity rendering quality at 2K resolution even under exaggerated expressions and driving hairs reasonably with the motion of the head. Project page: https://liaozhanfeng.github.io/HHAvatar.
Index Terms:
Head Avatar, Gaussian Splatting, Novel View Synthesis
@@ 1 Introduction
High-fidelity 3D human head avatar modeling is of great significance in many fields, such as VR/AR, telepresence, digital human and film production. Although some traditional head avatars [1, 2, 3, 4] realize high-fidelity animation, they typically require accurate geometries reconstructed and tracked from dense multi-view videos, thus limiting their applications in lightweight settings. On the other hand, recent works [5, 6, 7] have verified that Neural Radiance Fields (NeRF) [8] can skip the geometry reconstruction and the tracking steps but directly learn high-quality NeRF-based head avatars in dense or sparse views, greatly lowering the threshold for head avatar reconstruction. However, it still remains challenging for these NeRF-based approaches to synthesize high-fidelity images at 2K resolutions with pixel-level details, including hairs, wrinkles, and eyes.
Recently, 3D Gaussian Splatting (3DGS) [9], an explicit and efficient point-based representation, has been proposed for both high-fidelity rendering quality and real-time rendering speed. Compared to NeRF, the reconstruction quality of static and dynamic scenes [10, 11, 12] is much better while rendering time cost has been significantly reduced. Some works [13, 14, 15, 16, 17, 18, 19] have verified 3DGS can also create photorealistic head avatars that are controllable in terms of expression and pose. Nevertheless, these methods continue to face challenges in generating high-fidelity images at 2K resolutions with precise pixel-level details, faithfully representing highly complex and exaggerated facial expressions, and capturing dynamic hairs. To overcome this bottleneck and further improve the avatar quality, we propose Gaussian Head Avatar with Dynamic Hairs (HHAvatar), a novel representation that utilizes 3DGS for ultra high-fidelity head avatar with dynamic hair modeling.
@@
Previous explicit [20] and implicit [21, 22, 23] head avatars usually formulate the facial deformation via linear blend skinning (LBS) using the skinning weights and blendshapes like the FLAME model [24]. However, such a LBS-based formulation fails to represent exaggerated and fine-grained expressions by simple linear operations, limiting the representation ability of the head avatars. Inspired by NeRSemble [25], we propose a fully learnable expression-conditioned deformation field for the 3D head Gaussians, avoiding the limited capability of the LBS-based formulation. Specifically, we input the positions of the 3D Gaussians with expression coefficients into a MLP to directly predict the displacements from the neutral expression to the target one. Similarly, we control the motion of non-face areas, such as the neck, using the head pose as the condition. 3D Gaussian-based representation has the powerful ability to reconstruct high-frequency details, enabling our method to learn accurate deformation fields. In turn, the learned accurate deformation field facilitates the dynamic Gaussian head model to fit more dynamic details. As a result, our method is able to reconstruct finer-grained dynamic details of expressive human heads.
Figure 1: HHAvatar achieves ultra high-fidelity image synthesis with controllable expressions at 2K resolution. The above shows different identities animated by the same expression. The bottom shows that variations in hair positions can arise for identical poses, stemming from diverse hair status (i.e., position and speed) at the previous moment.
Unfortunately, even though using MLPs can fit more dynamic details, it is not easy to use the same method to model the physical motion of hairs. The reason is that the face can be determined by facial expressions and pose, but hair also needs to consider the speed and previous position due to inertia. Meanwhile, some works [26, 27, 28, 29, 30] mainly focus on hair reconstruction, but can not animate the head and the hair by facial expressions and head pose (i.e., not avatars). To this end, we propose to use more suitable representations for different head components to develop head avatars capable of simulating the physical motion of hair. Specifically, for hair modeling, we first obtain the 3DGS representation of canonical space, and then use MLP which takes hair status (i.e., position and speed) at the previous moment and the motion of the head as additional condition inputs, similar to deformable 3DGS [31], to deform and capture hair dynamics. Meanwhile, by employing mask supervision that introduces facial and hair occlusion, we segment the hair from the head and maintain the completeness of the head. Furthermore, we introduce a temporal module for fusing time series information as the motion of hair is influenced by the speed and position of the previous moment. As a result, our model can produce realistic animation with physical motion of the hairs.
@@
As a discretized representation, the gradients back-propagated to the 3D Gaussians cannot spread through the whole space. Thus the convergence of training heavily relies on a plausible initialization for both the geometry and the deformation field. However, simply initializing the 3D head Gaussians with a morphable template like FLAME [24] fails to model the long hairstyle and the shoulders. Hence, we further propose an efficient and well-designed geometry-guided initialization strategy. This strategy not only initializes the parts included in the morphable template, but also provides a better initialization for dynamic hairs. Specifically, instead of starting from stochastic Gaussians or a FLAME model, we initially optimize an implicit signed distance function (SDF) field along with a color field and a deformation MLP for modeling the basic geometry, color, and the expression-conditioned deformations of the head avatar respectively. The SDF field is converted to a mesh through Deep Marching Tetrahedra (DMTet) [32], with the color and deformation of the vertices predicted by the MLPs. Then we render the mesh and optimize them jointly under the supervision of multi-view RGB images. Finally, we use the mesh with per-vertex features from the SDF field to initialize the 3D Gaussians to lie on the basic head surface while the color and deformation MLPs are carried over to the next stage, ensuring stable training for convergence. The entire initialization process takes only around 10 minutes.
A preliminary version of this work has been published in CVPR 2024 [33], in which we propose a novel head avatar representation for modeling realistic animatable head avatars under lightweight sparse-view setups. In this paper, we extend it to handle avatar modeling of dynamic hairs. The contributions of our method can be summarized as:
@@
  • We propose HHAvatar, a new head avatar representation that employs controllable dynamic 3D Gaussians to model expressive human head avatars, producing ultra high-fidelity synthesized images at 2K resolutions. For modeling high-frequency dynamic details, we employ a fully learned deformation field upon the 3D head Gaussians, which accurately models extremely complex and exaggerated facial expressions and dynamic hair motion.
  • As far as we know, we propose the first head avatar that can model the physical motion of the hairs. By modeling hair separately from the head and introducing the kinematic features of the hair, the hair movement can be realistically simulated based on the motion of the head. By designing a training method that considers timing information and an occlusion perception module, dynamic hairs can be reconstructed under lightweight sparse-view setups.
@@
  • We design an efficient initialization strategy that leverages implicit representations to initialize the geometry and deformation, leading to efficient and robust convergence when training the HHAvatar. This strategy not only initializes the parts included in the morphable template, but also provides a better initialization for dynamic hairs.
@@
Benefiting from these contributions, our method surpasses recent state-of-the-art methods under lightweight sparse-view setups on the avatar quality by a large margin and model the physical motion of the hairs as shown in Fig. 1.
2 Related Work
In this section, we will briefly present the research related to head avatars and data-driven hair animation.
2.1 3D Head Avatar Modeling
@@ 2 Related Work
Due to the wide application value in the film and digital human industry, 3D head avatar reconstruction from multi-view images has always been a research hotspot. However, most of the works only focus on reconstruction without hair motion, and even only facial reconstruction. Traditional works [34, 35, 36, 37] reconstruct the scan geometry through multi-view stereo and then register a face mesh template to it. However, such methods usually require heavy computation. With the utilization of deep neural networks, current methods [38, 39, 40, 41] achieve very fast reconstruction, producing even more accurate geometry. Lombardi et al. [1], Bi et al. [42] and Ma et al. [3] represent the full head mesh through a deep neural network and train it with multi-view videos as supervision. However, due to the errors in geometric estimation, mesh-based head avatars typically suffer from texture blur. Therefore, some recent methods [43, 44] utilize NeRF representation [8] to synthesize novel view images without geometry reconstruction, or build NeRF on the head mesh template [2]. Furthermore, the NeRF-based methods are extended to sparse view reconstruction tasks [5, 7, 6, 25] and achieve impressive performance.
@@ 2.2 Head Avatar by Gaussian Splatting
Point elements as a discrete and unstructured representation can fit geometry with arbitrary topology [74] efficiently. Recent methods [75, 76, 77] open up a differentiable rasterization pipeline, such that the point-based representation is widely used in multi-view reconstruction tasks. Aliev et al. [78] and Ruckert et al. [79] propose to first render the feature map, which is transferred to the images through a convolutional renderer. Xu et al.[80] use neural point cloud associated with neural features to model a NeRF.
Recently, 3D Gaussian splatting [9] shows its superior performance, beating NeRF in both novel view synthesis quality and rendering speed. Some approaches [10, 12, 11, 81, 82, 83, 83, 84] extend Gaussian representation to dynamic scene reconstruction. However, these methods can not be migrated to the head avatar reconstruction tasks. GaussianAvatars [13] create photorealistic head avatars that are fully controllable in terms of expression and pose from multi-view videos. While other approaches like SplattingAvatar [14], PSAvatar [15], MonoGaussianAvatar [16], Rig3DGS [17], FlashAvatar [18] and GaussianHead [19] employ the coefficients of 3DMM to control the head and reconstruct a 3DGS-based animatable head avatar from monocular videos. However, these approaches need the hair maintain as still as possible and model the hair as a part rigidly attached to the head.
2.3 Data-driven Hair Animation
In academia and the film and gaming industries, using physics based simulations to create hair animations is a common practice [85]. However, using physics based simulations to generate hair animations may be computationally expensive. To address this issue, a simplified data-driven approach [86, 87, 88] simulates only a small portion of guided hair bundles and interpolates the remaining parts using skin weights learned from the complete simulation.
With the latest advances in deep learning, the use of neural networks has improved the efficiency of dynamic generation [89] and rendering [90, 91] of hair. Some methods [89] use deep neural networks for adaptive binding between normal hair and guided hair. Some methods [91] treat hair rendering as an image translation problem and generate realistic rendering of hair based on 2D hair masks and strokes. Some methods [90] achieve faster rendering and realistic results by using screen space neural rendering technology instead of the rendering part in the animation pipeline. However, these methods are still based on traditional hair simulation pipelines and use synthetic wigs, which require artists to manually set and are not easy to measure and evaluate. Some methods [92] propose to use a secondary motion graph for hair animation at runtime, without relying on traditional hair simulation pipelines. However, this method is limited by the artist’s design of wigs and control over hair simulation parameters, and cannot be used to animate real hair motions.
Additionally, certain techniques require only multi-view videos and camera parameters as input without necessitating the manual design of wigs by artists [28, 30, 93]. Some methods [28, 30] do not need the artist’s design of wigs. NeuWigs [28] uses the mixture of volumetric primitives (MVP) to model hair and does not rely on a manual hair design. However, this method needs a lot of cameras to capture a multi-view video for reconstruction, and only considers the driving of the hair and cannot complete the driving of the entire head. GaussianHair [30] and Gaussian Haircut [93] use specially designed Gaussian representation to model hair, which can achieve high-quality hair reconstruction. However, this method requires first relying on multi-view images to complete static reconstruction, and then driving based on a conventional CG rendering engine, which is not data-driven. Moreover, this method cannot animate the head by facial expressions and head poses. Compared with previous methods, our approach can achieve data-driven head modeling, including the face and the hair, under lightweight sparse-view setups.
@@ 3 Method
Figure 2: The pipeline of the HHAvatar rendering and reconstruction. We first optimize the guidance model including a neutral mesh, a deformation MLP and a color MLP in the Initialization stage. Then we use them to initialize the neutral Gaussians and the dynamic generator. Finally, 2K RGB images are synthesized through differentiable rendering and the super-resolution network, and the segmentation maps of the hair and the head are also synthesized through differentiable rendering. The HHAvatar are trained under the supervision of multi-view RGB videos and multi-view masks from face-parsing.
Figure 3: The detail of the temporal module. The input for the Hair Dynamic MLPs at time step $t$ is $\boldsymbol{X}_{0}$ (the position of the neutral Gaussian point), $\{X^{\prime}_{t-1},X^{\prime}_{t-2}\}$ (the position of the expressive Gaussian point at time step $t-1$ and time step $t-2$), and $\{\beta_{t},\beta_{t-1},\beta_{t-2}\}$ (the pose of the head at time step $t$, $t-1$, and $t-2$). The specific details of the Hair Dynamic MLPs are detailed in Sec. 3.1.2.
The pipeline of the reconstruction of HHAvatar is illustrated in Fig. 2, including the initialization stage and the training stage of HHAvatar. The detail of the temporal module is illustrated in Fig. 3. Before the beginning of the pipeline, we remove the background [94] of each image and jointly estimate the 3DMM model [65], 3D facial landmarks and the expression coefficients for each frame, and we obtain the mask of the hair and the head by face-parsing [95]  [^1] https://github.com/zllrunning/face-parsing.PyTorch. The whole model is optimized under the supervision of multi-view RGB videos and hair masks.
@@ 3.1 Avatar Representation
Our task is to reconstruct a dynamic head avatar controlled by facial expression, head pose, and head speed. For the region outside the hair on the head, the head avatar only depends on the facial expression and the head pose. For the hairs, a point $P_{t}$ in hairs at time step $t$ depend on the point $P_{t-1}$ at time step $t-1$, the speed of the point, the head pose at time step $t$, and the head speed. Therefore, we formulate the head avatar as dynamic 3D Gaussians conditioned on the facial expression, the head pose, head speed, and the points in hairs at the previous 2 time steps. To handle the dynamic changes, we input the above conditions to the head avatar model and output the position and other attributes of the Gaussians as above.
@@
Specifically, we first extract the neutral mesh for the head and the hair through DMTet [32] to initialize the neutral Gaussians (Sec. 3.3) and construct a canonical neutral Gaussian model with expression-independent attributes including $\boldsymbol{X}_{0}\in\mathbb{R}^{N\times 3}$ denotes the positions of the Gaussians with a neutral expression in the canonical space, $\boldsymbol{F}_{0}\in\mathbb{R}^{N\times 128}$ denotes the point-wise feature vectors as their intrinsic properties for generating high-resolution images, $\boldsymbol{Q}_{0}\in\mathbb{R}^{N\times 4}$, $\boldsymbol{S}_{0}\in\mathbb{R}^{N\times 3}$, and $\boldsymbol{A}_{0}\in\mathbb{R}^{N\times 1}$ denote the neutral rotation, scale and opacity respectively. $\boldsymbol{X}_{0}$, $\boldsymbol{F}_{0}$, $\boldsymbol{Q}_{0}$, $\boldsymbol{S}_{0}$, and $\boldsymbol{A}_{0}$ are fully optimizable. Note that we do not define the neutral color $\boldsymbol{C}_{0}$, but directly predict expression-dependent dynamic color from the point-wise feature vectors $\boldsymbol{F}_{0}$. Then, we construct a MLP-based expression conditioned dynamic generator $\boldsymbol{\Phi}$ to generate all the extra dynamic changes to the neutral model. Because the deformation of hair at time step $t$ is related to the hair position $X_{t-1}$ at time step $t-1$, the hair speed which can get from $X_{t-1}-X_{t-2}$, the head pose $\beta_{t}$ at time step $t$, and the head movement speed which can get from $\beta_{t}$ and $\beta_{t-1}$, the dynamic generator $\boldsymbol{\Phi}$ needs $X_{t-1}$, $X_{t-2}$, and $\beta_{t-1}$ as extra inputs. Overall, the whole HHAvatar can be formulated as:
$\begin{split}\{X_{t},C_{t},Q_{t},S_{t},A_{t}\}=\boldsymbol{\Phi}(\boldsymbol{X}_{0},\boldsymbol{F}_{0},\boldsymbol{Q}_{0},\boldsymbol{S}_{0},\boldsymbol{A}_{0};\\ X_{t-1},X_{t-2},\theta_{t},\beta_{t},\beta_{t-1}),\end{split}$ (2)
@@
with $\{X_{t},C_{t},Q_{t},S_{t},A_{t}\}$ denoting $\{X,C,Q,S,A\}$ at the time step $t$ and $\theta_{t}$ denoting expression coefficients at the time step $t$. During the training, we optimize all the parameters of the dynamic generator $\boldsymbol{\Phi}$ and the neutral Gaussian model $\{\boldsymbol{X}_{0},\boldsymbol{F}_{0},\boldsymbol{Q}_{0},\boldsymbol{S}_{0},\boldsymbol{A}_{0}\}$, which are highlighted in bold in the following.
3.1.1 Head Avatar